OpenStudy (anonymous):

You have 2m^3 of a mineral ore sample. It has a mass of 1 x 10^4 kg. Calculate the density of the ore sample and express your results in g/cm^3 with appropriate scientific notation. Please help I'm clueless on this one

OpenStudy (anonymous):

First, convert your starting quantities into the correct units, which are gram-mes and cubic centimeters. 10^4 kg = 10^7 g 2 cubic meters = 2 x 10^6 cubic centimeters Now divide the mass by the volume: 10^7/2 x 10^6 = 5 g/cm^3 hope that helped

OpenStudy (anonymous):

Thank you!! @warood

OpenStudy (anonymous):

your welcome!